A car stopped at a traffic light. It then travels along a straight road so that its distance from the light is given by xt=bt2-ct3 where, b=2.40m/s2 and c=0.120m/s3


A.) Calculate the average velocity of the car for the time interval t=0s to t=10.0s


B.) Calculate the instantaneous velocity of the car at t=0s; t=5.0s; and t=10.0s



1
Expert's answer
2021-03-30T06:44:06-0400

Answer

A. distance from the light is given by "x(t) =bt^2-ct^3"

where, b="2.40m\/s^2" and c="0.120m\/s^3"

distance from the light is given by

"x(t) =bt^2-ct^3" at t=0

x=0 m

At t=10sec

"x(t) =2.40\times 100-0.12\times1000"

=120m

the average velocity of the car for the time interval t=0s to t=10.0s

"V_a=\frac{dx}{dt}\\=\frac{120-0}{10-0}\\=12m/s"

B. the instantaneous velocity of the car at t=0s; t=5.0s; and t=10.0s

"v=\frac{dx}{dt}=2bt-3ct^2"

So at t=0 v=0

At t=5sec

"v=2\times2.4\times5-3\times0.12\times25\\=15m/s"

At t=10

"v=2\times2.4\times10-3\times0.12\times100\\=12m/s"